# Nonlinear Bayesian Filtering with Natural Gradient Gaussian Approximation

### v1

@@ V-B Stability Analysis
The stability analysis of our proposed NANO filter is based on the state-space model (1). We focus on the case where the process and measurement noises are both zero-mean Gaussian noise, satisfying $\mathbf{\xi}_{t}\sim\mathcal{N}(0,Q_{t})$ and $\mathbf{\zeta}_{t}\sim\mathcal{\mathcal{missing}}{N}(0,R_{t})$. The measurement-dependent loss $\ell(x_{t},y_{t})$ in the analysis is chosen to be the log-likelihood loss function that satisfies
$\ell(x_{t},y_{t})$ $=-\log p(y_{t}|x_{t})$
$=C_{0}+\frac{1}{2}(y_{t}-g(x_{t}))^{\top}R_{t}^{-1}(y_{t}-g(x_{t})).$
Here, $C_{0}>0$ is a constant value irrelevant to the state and measurement. For the purpose of stability analysis, we make the following regularity assumptions on the system functions $f$ and $g$.
@@
with $\lambda=\lambda_{\min}(P_{t|t-1}C^{\top}R_{t}^{-1}C)>0$ and $\lambda^{\prime}=\lambda_{\min}((F_{t-1}P_{t-1|t-1}F_{t-1}^{\top})^{-1}Q_{t})>0$, therefore (41) holds for some $\lambda_{1}>0$ for linear systems. Because of the continuity of (41) with respect to the observation function $g$, (41) also holds for those systems with almost linear measurement functions.
@@
Because $\bar{h}_{t}\equiv 0$ for linear observation functions, the constant $\delta_{1}$ above can be chosen to be sufficiently small for almost linear observation functions.
@@
With the fact that $\underline{p}I\leq\tilde{P}_{t|t}\leq\bar{p}I$, there exist $\epsilon_{1},\epsilon_{2}>0$, such that
$\mathbb{E}\left\{\left.(\|e_{t|t}\|-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(\|e_{t|t}\|-\epsilon_{1})\right|e_{t-1|t-1}\right\}$
$\leq$ $\frac{1}{1+\lambda_{0}}(\|e_{t-1|t-1}\|-\epsilon_{1})^{\top}\tilde{P}_{t-1|t-1}^{-1}(\|e_{t-1|t-1}\|-\epsilon_{1})+\epsilon_{2},$
@@
Also, if there exists $\tilde{\eta}>0$, such that $\tilde{\eta}<\|e_{t-1|t-1}\|<\eta$, for sufficiently small $\epsilon_{1},\epsilon_{2}$, we have the following supermartingale-like property:
$\mathbb{E}\left\{\left.(\|e_{t|t}\|-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(\|e_{t|t}\|-\epsilon_{1})\right|e_{t-1|t-1}\right\}$ (45)
$-(\|e_{t-1|t-1}\|-\epsilon_{1})^{\top}\tilde{P}_{t-1|t-1}^{-1}(\|e_{t-1|t-1}\|-\epsilon_{1})$
@@
Hence, if the initial estimation error $\|e_{0|0}\|<\eta$, then we can recursively use (V-B) to compute the estimation error at time $t$,
$\mathbb{E}\left\{(\|e_{t|t}\|-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(\|e_{t|t}\|-\epsilon_{1})\right\}$
@@
$\frac{1}{\bar{p}}\|\|e_{t|t}\|-\epsilon_{1}\|^{2}\leq(\|e_{t|t}\|-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(\|e_{t|t}\|-\epsilon_{1}),$
@@
$\|e_{t|t}\|^{2}\leq\left(\sqrt{\bar{p}(\|e_{t|t}\|-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(\|e_{t|t}\|-\epsilon_{1})}+\epsilon_{1}\right)^{2},$
for sufficiently small $\epsilon_{1}$, we obtained the exponentially bounded in mean square for $e_{t|t}$, i.e., there exist $\epsilon,\epsilon^{\prime}>0$, such that
@@ VI Robustifying NANO Filter with Gibbs Posterior
As illustrated in (6), the objective of update step of Bayesian filtering is to balance the information provided by the measurement data with that of the prior distribution. Nevertheless, acquiring the information from the measurement data necessitates a complete probabilistic model of the measurement-generating process. In other words, a perfect model of the output distribution $p(y_{t}|x_{t})$ must be known in advance. Unfortunately, specifying the true model of $p(y_{t}|x_{t})$ is often impossible due to sensor malfunctions or overlooked dynamics [35, 36], which introduces outliers in the measurement data. In this case, a viable solution is to replace the likelihood function term in (6) to a general measurement-dependent risk function, which leads to the following optimization problem:
@@
Here, $\ell^{G}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is the generalized measurement-dependent loss function. It turns out that the solution to the optimization problem (46) is known as the Gibbs posterior, which has an analytical form akin to the standard Bayesian posterior:
$p_{\mathrm{G}}(x_{t}|y_{1:t})=\frac{\exp\{\ell^{G}(x_{t},y_{t})\}p(x_{t}|y_{1:t-1})}{\int\exp\{\ell^{G}(x_{t},y_{t})\}p(x_{t}|y_{1:t-1})\,\mathrm{d}x_{t}}.$
@@
A notable observation is that the analysis of the Gaussian approximation in Section III-C and the derivation of the NANO filter in Section IV do not depend on the specific form of the loss function $\ell(x_{t},y_{t})$. Therefore, since Gibbs Bayesian filtering in (46) only modifies the loss function $\ell(x_{t},y_{t})$ in (6) to a more general form, $\ell^{G}(x_{t},y_{t})$, all of the previous analysis and algorithmic design can naturally extend to this case.
@@
$=$ $\begin{cases}-\log{\mathcal{N}(y_{t};h(x_{t}),1)}&\text{if}\quad|y_{t}-h(x_{t})|\leq\delta,\\ -\log{\mathrm{Laplace}(y_{t};h(x_{t}),\frac{1}{\delta})}&\text{otherwise.}\end{cases}$
=$\stackrel{{\scriptstyle c}}{{=}}$$\begin{cases}\frac{1}{2}|y_{t}-h(x_{t})|^{2}&\text{if}\quad|y_{t}-h(x_{t})|\leq\delta,\\ \delta\left(|y_{t}-h(x_{t})|-\frac{1}{2}\delta\right)&\text{otherwise.}\end{cases}$
Here, the notation ``=$\stackrel{{\scriptstyle c}}{{=}}"$ represents that two expressions are equivalent up to an additive constant, and $\delta>0$ is the threshold variable. In fact, $\ell^{\mathrm{h}}(x_{t},y_{t})$ represents the well-known Huber loss [37] used in robust regression, which is known to be less sensitive to outliers in data than the squared error loss. In practical applications, we utilize a smooth approximation of the Huber loss, known as the Pseudo-Huber loss, to improve optimization efficiency and ensure differentiability. The Pseudo-Huber loss is defined as:
$\ell^{\mathrm{ph}}(x_{t},y_{t})=\delta^{2}\left(\sqrt{1+\left(y_{t}-h(x_{t})\right)^{2}/{\delta}^{2}}-1\right).$ (47)
@@
Choice 3: Divergence-Dependent Loss Function: The final option is to utilize robust divergence to fit the data. This approach is based on the fact that minimizing the negative log-likelihood essentially amounts to minimizing the KL divergence between the true data distribution $p_{\mathrm{true}}(y_{t})$, which includes outliers, and the assumed likelihood. More specifically, the minimization of the measurement-dependent loss in (46) can be derived as follows:
@@
Formally analyzing which loss function provides superior robustness is challenging, as each is grounded in different principles. A common approach is to apply Huber’s robust statistics theory [41], which assesses the robustness of an estimator using its influence function. An estimator is considered robust if its influence function remains bounded as an outlier’s value increases indefinitely. Previous studies have conducted case-specific analyses of robustness for Huber loss [42], weighted loss [39], and $\beta$ loss [35] in linear Gaussian systems, demonstrating that adjusting the loss function can improve robustness. Since this paper focuses on applying robust loss functions to enhance the NANO filter, a detailed formal analysis is left for future work.
@@ VII Discussions
We utilize the natural gradient method to find the optimal Gaussian approximation for Bayesian filtering in this paper. Theoretically, the advantage of incorporating the natural gradient in Gaussian filtering may arise from the profound geometric properties of the statistical manifolds generated by the family of Gaussian distributions, as discussed in [50, 51]. Intuitively, the computation of the natural gradient involves second-order derivatives of the probability density functions. Thus, geometric concepts such as Riemannian metrics, curvatures, and geodesics on the statistical manifold can be engaged and reflected in the algorithm, leading to better numerical performance. Nevertheless, a precise description of the relationship between the natural gradient and Gaussian distribution remains an important research direction for further exploration.
Gradient-Based Gaussian Filters: Gradient-based Gaussian filters, which use gradient descent and its variants to solve Gaussian filtering, were initially proposed in [52], utilizing Monte Carlo techniques to construct gradient descent information. Subsequently, cubature rules were introduced to replace the Monte Carlo method, and various optimization approaches, such as the conditional gradient method and the alternating direction method of multipliers, were employed to replace basic gradient descent [53, 54, 55].
@@
Unfortunately, these methods have certain assumptions and limitations compared to our approach. For instance, (i) the methods in [52, 53, 54] are restricted to systems with Gaussian noise; (ii) [52] and [55] require performing linear approximations of the measurement model; and (iii) [53] modifies the implicit hard constraints of the original problem into soft constraints, rendering it inequivalent to the original problem. Also, they both lack rigorous proofs of algorithm convergence and stability guarantees for the estimation error, which we provide in Section V. Additionally, their optimization problems are constructed based on direct minimization of forward KL divergence, which does not support the extension to the Gibbs posterior as proposed in our method.
@@
Also, we would like to emphasize that, although these works employ gradient methods for Gaussian approximation, they do not theoretically analyze the stationary points corresponding to the optimal Gaussian approximation. For example, they do not discuss the rationale for using gradient-based methods in the update process, nor do they consider the Gaussian approximation in the prediction step. In contrast, our paper systematically discusses both the prediction and update steps, providing a comprehensive analysis of the stationary points for the optimal Gaussian approximation.
@@
Kalman Filtering as Natural Gradient Descent: Corollary 1 shows that the canonical Kalman Filter (KF) can be interpreted as a single iteration of the natural gradient in our proposed NANO filter. This result is similar to the findings in [56], where the equivalence between the KF and online natural gradient descent is established. However, [56] focuses on parameter estimation, where the state is treated as a deterministic variable representing the parameters of the measurement model. In that context, the natural gradient is applied to maximize the likelihood with respect to the measurement model. In contrast, our work considers the more general linear Gaussian filtering problem, where the state is a random variable, and the natural gradient is used to update the cost with respect to the posterior distribution of the state. Therefore, our result is a generalization of the result in [56].
@@ VIII Experiments
In this section, we validate the proposed NANO filter and its robust variants using the loss functions introduced in Section VI. We conduct simulations for both linear and nonlinear systems, along with a real-world experiment. All the simulations and experiment are evaluated by root mean square error which is defined by
@@ VIII-A Linear System Simulation: Wiener Velocity Model
First, we perform simulations for the canonical linear Gaussian system, specifically using the Wiener velocity model that is commonly employed for target tracking simulations [10]. In this model, the state represents the position and velocity of a moving object in two dimensions. The state vector is defined as $x=\left[p_{x}\ p_{y}\ v_{x}\ v_{y}\right]^{\top}$, where $p_{x}$ and $p_{y}$ are the object’s positions along the longitude and lateral directions, and $v_{x}$ and $v_{y}$ are the corresponding velocities. The measurements are direct, noisy observations of the position components. The state transition model with a time step $\Delta t=0.1$ and the measurement model can be described as
@@
Figure 1: Box plot of RMSE of KF, UKF, PLF and NANO filter, for the standard Wiener velocity model.
Figure 2: Box plot of RMSE for KF, UKF, PLF, NANO filter, and its robust variants: $\beta$-NANO filter, Huber-NANO filter, and Weight-NANO filter with various parameter values, for the Wiener velocity model with measurement outliers. Note that the small black square “
@@
To validate our method, we consider two scenarios. In the first scenario, the proposed NANO filter is compared against the KF, the optimal filter for linear Gaussian systems, as well as UKF and PLF. As shown in Fig. 1, the NANO filter achieves the same performance as the KF family within one iteration, supporting the result in Corollary 1 that KF equals a single iteration of NANO filter. In the second scenario, we consider the case where the measurement data deviates from the measurement model due to contamination of outliers. Specifically, the measurement data has a 0.1 probability of being contaminated, modeled as $\zeta_{t}\sim 0.9\cdot\mathcal{N}(0,R)+0.1\cdot\mathcal{N}(0,1000R)$. For this case, as illustrated in Fig. 2, we evaluate the robust variants of the NANO filter, including the Huber-NANO filter, Weight-NANO filter, and $\beta$-NANO filter, which utilize the robust loss functions $\ell^{\mathrm{ph}}$ (47), $\ell^{\mathrm{w}}$ (48), and $\ell^{\beta}$ (50), respectively. The box plots indicate that the robust variants of the NANO filter generally outperform the standard NANO filter using different parameters. Notably, the Weight-NANO filter with $c=25$ shows the best performance, significantly improving estimation accuracy.

### v2

@@ V-B Stability Analysis
Next, we will analyze the stability of the proposed NANO filter. Stability is the most critical property, as it ensures that the estimation error remains bounded throughout the filtering process. Our stability analysis is conducted based on the state-space model given in (1). We focus on the case where the process and measurement noise are both zero-mean Gaussian noise, satisfying $\mathbf{\xi}_{t}\sim\mathcal{N}(0,Q_{t})$ and $\mathbf{\zeta}_{t}\sim\mathcal{\mathcal{missing}}{N}(0,R_{t})$. In this case, the measurement-dependent loss $\ell(x_{t},y_{t})$ is the log-likelihood loss function that satisfies $\ell(x_{t},y_{t})=-\log p(y_{t}|x_{t})=C_{0}+\frac{1}{2}(y_{t}-g(x_{t}))^{\top}R_{t}^{-1}(y_{t}-g(x_{t}))$. Here, $C_{0}>0$ is a constant value irrelevant to the state and measurement. For the purpose of stability analysis, we make the following regularity assumptions on the system functions $f$ and $g$.
@@
with $\lambda=\lambda_{\min}(P_{t|t-1}C^{\top}R_{t}^{-1}C)>0$ and $\lambda^{\prime}=\lambda_{\min}((F_{t-1}P_{t-1|t-1}F_{t-1}^{\top})^{-1}Q_{t})>0$, therefore (41) holds for some $\lambda_{1}>0$ for linear systems. Because of the continuity of (41) with respect to the measurement function $g$, (41) also holds for those systems with almost linear measurement functions.
@@
Because $\bar{h}_{t}\equiv 0$ for linear measurement functions, the constant $\delta_{1}$ above can be chosen to be sufficiently small for almost linear measurement functions.
@@
With the fact that $\underline{p}I\leq\tilde{P}_{t|t}\leq\bar{p}I$, there exist a vector $\epsilon_{1}\in\mathbb{R}^{n}$ and a constant $\epsilon_{2}>0$, such that
$\mathbb{E}\left\{\left.(e_{t|t}-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(e_{t|t}-\epsilon_{1})\right|e_{t-1|t-1}\right\}$ (45)
$\leq$ $\frac{1}{1+\lambda_{0}}(e_{t-1|t-1}-\epsilon_{1})^{\top}\tilde{P}_{t-1|t-1}^{-1}(e_{t-1|t-1}-\epsilon_{1})+\epsilon_{2},$
@@
Also, if there exists $\tilde{\eta}>0$, such that $\tilde{\eta}<\|e_{t-1|t-1}\|<\eta$, for $\epsilon_{1}$ with sufficiently small norm and sufficiently small $\epsilon_{2}$, we have the following supermartingale-like property:
$\mathbb{E}\left\{\left.(e_{t|t}-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(e_{t|t}-\epsilon_{1})\right|e_{t-1|t-1}\right\}$
$-(e_{t-1|t-1}-\epsilon_{1})^{\top}\tilde{P}_{t-1|t-1}^{-1}(e_{t-1|t-1}-\epsilon_{1})$
@@
Hence, if the initial estimation error $\|e_{0|0}\|<\eta$, then we can recursively use (45) to compute the estimation error at time $t$,
$\mathbb{E}\left\{(e_{t|t}-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(e_{t|t}-\epsilon_{1})\right\}$
@@
$\frac{1}{\bar{p}}\|e_{t|t}-\epsilon_{1}\|^{2}\leq(e_{t|t}-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(e_{t|t}-\epsilon_{1}),$
@@
$\|e_{t|t}\|\leq\sqrt{\bar{p}(|e_{t|t}|-\epsilon_{1})^{\top}\tilde{P}_{t|t}^{-1}(|e_{t|t}|-\epsilon_{1})}+|\epsilon_{1}|,$
and we obtained the exponentially bounded in mean square for $e_{t|t}$, i.e., there exist $\epsilon,\epsilon^{\prime}>0$, such that
@@ VI Robustifying NANO Filter with Gibbs Posterior
In the update step of Bayesian filtering, as illustrated in (6), the objective is to strike a balance between the information provided by the measurement data and the prior distribution. However, obtaining reliable information from the measurement data requires a comprehensive probabilistic model of the measurement-generating process. In other words, a precise specification of the output distribution $p(y_{t}|x_{t})$ is needed beforehand. Unfortunately, in practical scenarios, it is often challenging to specify the true model of $p(y_{t}|x_{t})$ due to factors such as sensor malfunctions or unmodeled system dynamics. This leads to a mismatch between the true data-generating process and the assumed output probability model. Measurements influenced by such model misspecifications, which are typically referred to as outliers, require special attention to maintain the reliability of state estimates [35, 36]. One effective approach to handle this issue is to replace the likelihood function in (6) with a generalized measurement-dependent loss function, leading to the following variational problem:
@@
Here, $\ell^{G}:\mathbb{R}^{n}\times\mathbb{R}^{m}\to\mathbb{R}$ is the generalized measurement-dependent loss function. It turns out that the solution to (46) is known as the Gibbs posterior, which has an analytical form akin to the standard Bayesian posterior:
$p_{\mathrm{G}}(x_{t}|y_{1:t})=\frac{\exp\{-\ell^{G}(x_{t},y_{t})\}p(x_{t}|y_{1:t-1})}{\int\exp\{-\ell^{G}(x_{t},y_{t})\}p(x_{t}|y_{1:t-1})\,\mathrm{d}x_{t}}.$
@@
A notable fact is that the analysis of the Gaussian approximation in Section III-C and the derivation of the NANO filter in Section IV do not depend on the specific form of the loss function $\ell(x_{t},y_{t})$. Therefore, since Gibbs Bayesian filtering in (46) only modifies the loss function $\ell(x_{t},y_{t})$ in (6) to a more general form, $\ell^{G}(x_{t},y_{t})$, all of the previous analysis and algorithmic design can naturally extend to this case.
@@
$=$ $\begin{cases}-\log{\mathcal{N}(y_{t};g(x_{t}),1)}&\text{if}\quad|y_{t}-g(x_{t})|\leq\delta,\\ -\log{\mathrm{Laplace}(y_{t};g(x_{t}),\frac{1}{\delta})}&\text{otherwise.}\end{cases}$
=$\stackrel{{\scriptstyle c}}{{=}}$$\begin{cases}\frac{1}{2}|y_{t}-g(x_{t})|^{2}&\text{if}\quad|y_{t}-g(x_{t})|\leq\delta,\\ \delta\left(|y_{t}-g(x_{t})|-\frac{1}{2}\delta\right)&\text{otherwise.}\end{cases}$
Here, the notation ``=$\stackrel{{\scriptstyle c}}{{=}}"$ represents that two expressions are equivalent up to an additive constant, and $\delta>0$ is the threshold variable. In fact, $\ell^{\mathrm{h}}(x_{t},y_{t})$ represents the famous Huber loss [37] used in robust regression, which is known to be less sensitive to outliers in data than the squared error loss. In practical applications, we utilize a smooth approximation of the Huber loss, known as the Pseudo-Huber loss, to improve optimization efficiency and ensure differentiability. The Pseudo-Huber loss is defined as:
$\ell^{\mathrm{ph}}(x_{t},y_{t})=\delta^{2}\left(\sqrt{1+\left(y_{t}-g(x_{t})\right)^{2}/{\delta}^{2}}-1\right).$ (47)
@@
Choice 3: Divergence-Dependent Loss Function: The final option is to utilize robust divergence to fit the data. This approach is based on the fact that minimizing the negative log-likelihood essentially amounts to minimizing the KL divergence between the true data distribution $p_{\mathrm{true}}(y_{t})$, which includes outliers, and the assumed likelihood:
@@
Formally analyzing which loss function provides superior robustness is challenging as each is grounded in different principles. A common approach is to apply Huber’s robust statistics theory [40], which assesses the robustness of an estimator using its influence function. An estimator is considered robust if its influence function remains bounded as an outlier’s value increases indefinitely. Previous studies have conducted case-specific analyses of robustness for Huber loss [41], weighted loss [38], and $\beta$ loss [35] in linear Gaussian systems, demonstrating that adjusting the loss function can improve robustness. Since this paper focuses on applying robust loss functions to enhance the NANO filter, a detailed formal analysis is left for future work.
@@ VII Discussions
We utilize the natural gradient method to find the optimal Gaussian approximation for Bayesian filtering in this paper. The advantage of incorporating the natural gradient in Gaussian filtering may arise from the profound geometric properties of the statistical manifolds generated by the family of Gaussian distributions, as discussed in [49, 50]. Intuitively, the computation of the natural gradient involves second-order derivatives of the probability density functions. Thus, geometric concepts such as Riemannian metrics, curvatures, and geodesics on the statistical manifold can be engaged and reflected in the algorithm, leading to better convergence performance. Nevertheless, a precise description of the relationship between the natural gradient and Gaussian distribution remains an important research direction for further exploration.
@@
Gradient-Based Gaussian Filters: Gradient-based Gaussian filters, which use gradient descent and its variants to solve Gaussian filtering, were initially proposed in [51], utilizing Monte Carlo techniques to approximate the exact gradient. Subsequently, cubature rules were introduced to replace the Monte Carlo method for gradient approximation, and various optimization approaches, such as the conditional gradient method, the alternating direction method of multipliers, and the natural gradient descent were employed to replace basic gradient descent for optimization [52, 53, 54, 55]. These methods have certain assumptions and limitations compared to our approach. For instance, (i) the methods in [51, 52, 53, 55] are restricted to systems with Gaussian noise; (ii) [51, 54] and [55] require performing linear approximations of the measurement model; (iii) [52] modifies the implicit hard constraints of the original problem into soft constraints, rendering it inequivalent to the original problem and (iv) [56] only finds the maximum a posterior estimate, rather than the entire optimal Gaussian distribution. Also, they both lack rigorous proofs of algorithm convergence and stability guarantees for the estimation error, which we provide in Section V. Besides, their optimization problems are constructed based on direct minimization of forward KL divergence between the candidate distribution and the true posterior, which does not support the extension to the Gibbs posterior as proposed in our method.
@@
Moreover, we would like to emphasize that, while these works utilize gradient-based methods for Gaussian approximation, they lack a theoretical analysis of the stationary points related to the optimal Gaussian approximation. Specifically, they neither justify the use of gradient methods in the update process nor address the Gaussian approximation in the prediction step. In contrast, our paper provides a comprehensive analysis covering both the prediction and update steps, including a detailed examination of the stationary points for the optimal Gaussian approximation.
@@
Kalman Filtering as Natural Gradient Descent: Corollary 1 shows that the canonical KF can be interpreted as a single iteration of the natural gradient in our proposed NANO filter. This result looks similar to the findings in [57], where the equivalence between KF and online natural gradient descent is established. However, [57] focuses on parameter estimation, where the state is treated as a deterministic variable representing the parameters of the measurement model. In that context, the natural gradient is applied to maximize the likelihood with respect to the measurement model. In contrast, our result is built on a more general Bayeisan view, where the state is a random variable, and the natural gradient is used to find the optimal parameters of state distribution. Therefore, our result can be treated as a generalization of the result in [57].
@@ VIII Experiments
In this section, we validate the proposed NANO filter and its robust variants using the loss functions introduced in Section VI. We conduct simulations for both linear and nonlinear systems, along with a real-world experiment. All the simulations and experiment are evaluated by root mean square error (RMSE) which is defined by
@@ VIII-A Linear System Simulation: Wiener Velocity Model
First, we perform simulations for Wiener velocity model, which is a canonical linear Gaussian model commonly employed for target tracking [10]. In this model, the state represents the position and velocity of a moving object in two dimensions. The state vector is defined as $x=\left[p_{x}\ p_{y}\ v_{x}\ v_{y}\right]^{\top}$, where $p_{x}$ and $p_{y}$ are the object’s positions along the longitude and lateral directions, and $v_{x}$ and $v_{y}$ are the corresponding velocities. The measurements are direct, noisy observations of the position components. The state transition model with a time step $\Delta t=0.1$ and the measurement model can be described as
@@
Figure 1: Box plot of RMSE of KF, UKF, PLF and NANO filter, for the standard Wiener velocity model. Note that the small black square “
@@
Figure 2: Box plot of RMSE for KF, UKF, PLF, NANO filter, and its robust variants with various parameter values, for the Wiener velocity model with measurement outliers.
To validate our method, we consider two scenarios. In the first scenario, the proposed NANO filter is compared against the KF, the optimal filter for linear Gaussian systems, as well as UKF and PLF. As shown in Fig. 1, the NANO filter achieves the same performance as the KF family within one iteration, supporting the result in Corollary 1 that KF equals a single iteration of NANO filter. In the second scenario, we consider the case where the measurement data deviates from the measurement model due to contamination of outliers. Specifically, the measurement data has a 10% probability of being contaminated, modeled as $\zeta_{t}\sim 0.9\cdot\mathcal{N}(0,R)+0.1\cdot\mathcal{N}(0,1000R)$. For this case, as illustrated in Fig. 2, we evaluate the robust variants of the NANO filter, including the Huber-NANO filter, Weight-NANO filter, and $\beta$-NANO filter, which utilize the robust loss functions $\ell^{\mathrm{ph}}$ (47), $\ell^{\mathrm{w}}$ (48), and $\ell^{\beta}$ (50), respectively. The box plots indicate that the robust variants of the NANO filter generally outperform the standard NANO filter using different parameters. Notably, the Weight-NANO filter with $c=25$ shows the best performance.